OpenStudy (anonymous):

find the equation of a parabola whose vertex is on the y axis, latus rectum 5, passing through (-3/2, -16/5) and whose axis on the y-axis. The answers are x^2 = 5(y+ 73/20) and x^2 = -5(y+11/4). HELP ME PLEASE. Show it to me step by step.

OpenStudy (anonymous):

kk we can do this

OpenStudy (anonymous):

we know the \(x\) coordinate of the vertex is \(0\) because it is on the \(y\) axis

OpenStudy (anonymous):

and we know \(4p=5\) or \(4p=-5\) because of that "latus rectum" bit

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

since it has axis of symmetry as the \(y\) axis you know the \(x\) part is squared

OpenStudy (anonymous):

so it looks like \[x^2=5(y-k)\] or \[x^2=-5(y-k)\]

OpenStudy (anonymous):

put \(x=-\frac{3}{2}\) and \(y=-\frac{16}{5}\) and solve for \(k\)

OpenStudy (anonymous):

i am going to go to bed soon, and try to get this "rectum" business out of my head

OpenStudy (anonymous):

yes! Thank you so much :)) Goodnight!

OpenStudy (anonymous):

wait. how bout 73/20? the other one?

OpenStudy (anonymous):

okay got it now

OpenStudy (anonymous):

you have two equations to solve for \(k\) one with \(5\) and the other with \(-5\) i didn't do it, but i imagine that is how you are going to get two answers

OpenStudy (anonymous):

thank you :)))

OpenStudy (anonymous):

yw